Find the (i) resonant and (ii) anti-resonant frequencies for series and parallel LCR circuits, respectively with the following data: L=10 mH, C=10 μF and R=1 kΩ. 


1
Expert's answer
2022-03-17T09:30:41-0400

Answer

The resonant frequency is given

"w=\sqrt{\frac{1}{10*10^{-3}*10*10^{-6}}}\\=3.16*10^3rad/sec"

Frequecy

"f=\frac{1}{2\pi \sqrt{LC}}\\=0.50*10^3Hz"